# Safe Learning for Contact-Rich Robot Tasks: A Survey from Classical Learning-Based Methods to Safe Foundation Models

### v1

@@ 3.1.1 Safe Exploration (safe learning before contact)
A major research direction in safe execution for contact-rich tasks is the use of RL to automatically tune the parameters of compliant controllers [elguea2023review], such as impedance or admittance control. In these approaches, RL algorithms are employed to learn optimal values for parameters such as stiffness [zhang2024srl], damping, and inertia, enabling the robot to adapt its compliance to varying task requirements and contact conditions [martinez2015safe]. This data-driven adaptation helps maintain safety by minimizing excessive contact forces, preventing instability, and improving task performance.
@@
Recent research has demonstrated the effectiveness of RL-based parameter learning in a variety of contact-rich scenarios. For example, RL has been used to adjust impedance parameters online to ensure safe and robust assembly[stan2020reinforcement, 192, 49, 174], insertion [55, 279, 344], and surface interaction tasks [61]. Some works incorporate safety constraints directly into the RL framework, such as limiting the maximum allowable force or ensuring stability margins during learning [287, 66, zhang2024srl]. Others combine RL with model-based safety filters [tang2024learning] or control barrier functions [zhang_enhancing_nodate] to guarantee constraint satisfaction throughout the learning and execution process. [zhang2024srl]
@@
Large robot foundation models enhance safe execution by converting unstructured scene and instruction inputs into structured, verifiable manipulation plans, then constraining low‑level interaction. Plan–parameterize–enforce (PPE). High-level language/vision modules parse tasks and safety rules (e.g., force caps, no-touch regions, human proximity) and compile them into machine-checkable constraints; mid-level policies parameterize compliant skills (pose/force/impedance) rather than direct torques; low-level safety filters (CBF/reachability/MPC shields) project commands online to maintain invariance and limits [330]. Constrained VLA (e.g., “safeVLA”, “TLA”-style) Semantic policies are restricted via schema-conformant action spaces and pre-execution checks; code-as-policies and structured prompting inject explicit guardrails (caps, forbidden zones) before motion is issued [33, 330, 10705419]. Semantic-to-physical constraint grounding. VLMs/VLAs translate natural-language rules and scene understanding (affordances, fragile surfaces) into explicit force/region/speed constraints that downstream safety layers can enforce [376].
@@ 3.1.3 Provable Safety Methods
Robotic safety has long been researched in classical control before learning-based methods became prevalent [hogan1985impedance, khalil2002nonlinear, bansal2017hamilton, hogan2022contact]. Rigorous mathematical tools were developed to provide provable guarantees of stability and safety [blanchini1999set, blanchini2008set, slotine1991applied, kumar_safe_2025]. This section aims to systematically describe foundational methods for certifying safety, including Lyapunov-based methods, control barrier functions (CBF), contraction metrics, reachability analysis and invariant sets. While originally developed for classical model-based control, many of these tools are increasingly integrated in modern learning-based frameworks [wabersich2023data]. This section aims to establish the control theoretical foundations for understanding and developing safe learning methods for contact-rich robotic tasks.
@@
Lyapunov-based methods build on the concept of a Lyapunov function, a scalar energy-like function $V(\boldsymbol{x})$ that decreases along trajectories of a dynamical system. Formally, consider a system $\dot{\boldsymbol{x}}={f}(\boldsymbol{x},\boldsymbol{u})$ with state $\boldsymbol{x}\in\mathcal{X}$ and control input $\boldsymbol{u}\in\mathcal{U}$. A continuously differentiable function $V:\mathbb{R}^{n}\to\mathbb{R}_{\geq 0}$ is a Lyapunov function for a closed-loop system if it is positive definite ($V(\boldsymbol{0})=0$, $V(\boldsymbol{x})>0$ for $\boldsymbol{x}\neq\boldsymbol{0}$) and its time derivative is negative semi-definite ($\dot{V}(\boldsymbol{x})\leq 0\,\forall x$). For safety, Lyapunov analysis is often used to certify invariant sets. If we can identify a set of safe states $\mathcal{S}(e.g.\,\mathcal{S}=\{\boldsymbol{x}\mid V(\boldsymbol{x})\leq c\})$ and a Lyapunov-like function $V(\boldsymbol{x})$ that decreases or stays constant on $\mathcal{S}$, then any trajectory starting in $\mathcal{S}$ will remain in $\mathcal{S}$ for all time, making $\mathcal{S}$ a forward-invariant safe set [khalil2002nonlinear].
@@
Lyapunov-based methods are well-suited to scenarios where the main objective is stability, for example, regulating contact force to a desired value. Nearly all classical robot controllers, e.g., PD controllers or compliance controllers, are derived with an implicit Lyapunov function to prove stability [16, 2, 32]. Passivity-based control is another Lyapunov-related approach. By designing the closed-loop system to be passive, one can guarantee stable interactions, which means the robot can safely interact with the environment because it will not generate energy that could cause instability [van2014port, forni2018differential, hill2022dissipativity].
@@
Despite their broad use, a major challenge for Lyapunov methods is finding a suitable Lyapunov function for complex nonlinear systems [khalil2002nonlinear]. For high-dimensional robots with complex contact dynamics, hand-designing a $V(\boldsymbol{x})$ that decreases in all conditions, especially across contact/non-contact modes, is difficult. In hybrid systems, standard Lyapunov theory does not directly apply and one may need multiple Lyapunov functions [branicky2002multiple, giesl2015review, manchester2017control]. Another limitation is that classical Lyapunov methods are primarily designed to certify stability or asymptotic convergence of the closed-loop system. Unless the Lyapunov sublevel sets are explicitly constructed to be forward invariant and contained in a prescribed safe set, Lyapunov stability alone does not guarantee that state or input constraints are respected along the entire transient [romdlony2016stabilization, li2023survey]. In particular, a Lyapunov-stable controller may drive a robot back to its desired equilibrium after a disturbance but still allow temporary violations of safe force or workspace limits [manjunath2021safe, wu2015safety]. Therefore, Lyapunov-based designs are well-suited for ensuring long-term stability or boundedness, but typically need to be augmented with barrier-type conditions to enforce state constraints and safety [ames2019cbf, blanchini1999set].
@@
Control barrier functions (CBFs) provide a formal, Lyapunov-like tool specifically for enforcing safety constraints. While Lyapunov functions certify the stability of an equilibrium, barrier functions certify forward invariance of a safe set, which prove the system will never enter certain unsafe regions [romdlony2016stabilization, ames2016control, ames2019cbf]. The typical setup is as follows: Define a continuously differentiable scalar boundary function $h(\boldsymbol{x})$ such that the safe set $\mathcal{C}$ is $\{\boldsymbol{x}\mid h(\boldsymbol{x})\geq 0\}$ and the unsafe set is $\{\boldsymbol{x}\mid h(\boldsymbol{x})<0\}$. As an example, the boundary $h(\boldsymbol{x})=0$ can describe the contact force reaching a limit. A valid CBF satisfies a condition that ensures $h(\boldsymbol{x})$ cannot decrease into the negative unsafe region. One common form: There exists an extended class-$\mathcal{K}$ function $\alpha(\cdot)$ such that for all states on the boundary of $\mathcal{C}$, we can choose control inputs $\boldsymbol{u}$ to yield $\dot{h}(\boldsymbol{x},\boldsymbol{u})\geq-\alpha(h(\boldsymbol{x}))$. If this holds, then whenever $h(\boldsymbol{x})$ is zero or positive, its derivative cannot drive $h(\boldsymbol{x})$ across the zero boundary into the unsafe region. Geometrically, $h(\boldsymbol{\boldsymbol{x}})$ can be seen as a ”barrier” that goes to zero at the safety boundary and typically to negative values outside the safe set. The CBF condition forces the controlled dynamics to hover at or increase $h(\boldsymbol{x})$ at the boundary, never allowing a drop below zero [wu2015safety].
@@
CBFs are especially powerful for real-time enforcement of safety constraints. Unlike open-loop potential field methods [batinovic2022path, pan2021improved], CBFs offer a guarantee of safety because the CBF condition is a hard constraint on the control input.
@@
Despite their strengths, CBFs come with several challenges. First, they require a reasonably accurate model of the system and environment. Model uncertainty or external disturbances can break the guarantee. There is active research on robust CBFs and adaptive CBFs to account for uncertainty, but guaranteeing safety under uncertainty often leads to conservative controllers [pati2023robust, buch2021robust, wang2025safe]. Second, there can be conflicts between multiple CBFs or between a CBF and control goals. It is possible that no single controller can simultaneously satisfy all constraints and achieve the desired motion. In such cases, typically one should relax constraints [isaly2024feasibility, breeden2023compositions]. Third, the design of CBFs becomes significantly more involved for safety constraints with high relative degree, such as contact and force constraints that depend on higher-order derivatives of the state, which are ubiquitous in contact-rich robotic tasks [breeden2021high, wang2022high, wong2025contact].
@@
Contraction metrics offer a fundamentally different viewpoint than the Lyapunov-based and CBF methods that provide scalar certificates for stability or safety with respect to particular equilibria or safe sets. Rather than analyzing the behavior of trajectories relative to a fixed point or a boundary, contraction theory studies the evolution of the differential distance between arbitrary trajectories of a dynamical system. A system is said to be contracting if infinitesimal displacements between trajectories converge exponentially to zero under an appropriate Riemannian metric. When such a contraction metric exists, all trajectories within a contraction region converge exponentially toward one another, independent of their initial conditions. This yields a strong form of incremental stability: not only do trajectories converge to a specific equilibrium, but they converge toward a unique flow or invariant trajectory structure of the system. The contraction metric thus serves as the formal certificate ensuring this incremental convergence and generalizes classical Lyapunov analysis to trajectory-to-trajectory stability [lohmiller1998contraction, slotine2003modular, manchester2017control].
@@
This condition means the distance between two nearby solutions measured by $\boldsymbol{M}$ decays at least at rate $\lambda$. One can interpret it as a generalized Lyapunov condition applied to the difference dynamics. If we consider the error $|\tilde{\boldsymbol{x}}|=\boldsymbol{x}_{1}-\boldsymbol{x}_{2}$ between two trajectories $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, contraction theory guarantees $|\tilde{\boldsymbol{x}}|\leq ke^{-\lambda t}|\tilde{\boldsymbol{x}}(\boldsymbol{0})|$ for some $k>0$. A remarkable property is that if a system is contracting, it has a unique equilibrium or invariant trajectory that is globally exponentially stable [lohmiller1998contraction]. In control contexts, we often speak of control contraction metrics (CCMs). A CCM extends the problem to controlled systems. It provides a condition under which a feedback controller can be designed so that the closed-loop system is contracting. If a CCM exists, one can synthesize a controller that achieves exponential tracking of any feasible reference trajectory [manchester2017control].
@@
Contraction metrics are particularly useful in scenarios requiring robust trajectory tracking and disturbance rejection. A major challenge of this approach is finding a contraction metric for a given system, which can be challenging as finding a Lyapunov function. For complicated robots, $\boldsymbol{M}(\boldsymbol{x})$ might need to be very complex or even non-existent in closed form [giesl2021computing, giesl2022review]. Another challenge is that contraction typically addresses stability properties but not explicit state constraints [singh2017robust].
Reachability analysis and invariant sets approaches compute the set of states from which the system can remain safe indefinitely or the set of states that could lead to an unsafe condition. Techniques like Hamilton-Jacobi (HJ) reachability solve a PDE to find the backward reachable set of an unsafe target [mitchell2005time, bansal2017hamilton]. Consider a system whose dynamics are described by an ordinary differential equation $\dot{\boldsymbol{x}}=\boldsymbol{f}(\boldsymbol{x}(t),\boldsymbol{u}(t))$ where $\boldsymbol{x}$ is the system’s states, $\boldsymbol{u}(t)$ is the control input. To compute the reachable set, a value function $V(\boldsymbol{x},t)$ is typically defined, representing whether the system can reach the target set from state $\boldsymbol{x}$ at time $t$. This value function satisfies the HJ equation:
@@
By solving this equation, the boundary of the reachable set can be obtained, allowing for the analysis of the system’s behavior under various control strategies [mballo2025hamilton, choi2023forward, liu2024hamilton].
The limitation is the curse of dimensionality. Solving HJ PDEs or exhaustive reachability is only feasible for simplified systems [darbon2016algorithms, bansal2021deepreach, ganai2023iterative]. In safe learning contexts, reachability is often used as a runtime safety filter [Shao2021RTS, selim2022safe, kochdumper2023provably].
@@ 3.2 Contact-Rich Tasks
Our categorization is application-independent, however, it is possible to categorize the tasks also w.r.t. the application areas such as industry, household and healthcare [Tsuji2025asu]. Many contact-rich tasks are demonstrated for household applications that cannot be automated, as opposed to industry. These include general household tasks such as wiping a table or a vase [151, 6], inserting a plug [112, 146], and opening a door [36, 149]; and kitchen tasks such as cutting vegetables [21, 157], stirring or scooping food [92, 191], and rolling pizza [156]. Industry applications usually focus on flexible manufacturing tasks such as inserting a gear [177, 192], nut threading [10], building furniture [43], and disassembly [20, 62]. Healthcare applications mostly include the surgery operations [339, 188] and physically assisting humans (Sec. 3.2.4).

### v2

@@ 3.1 Safe Learning for Contact-Rich Tasks
A major research direction in safe execution for contact-rich tasks is the use of RL to automatically tune the parameters of compliant controllers [70], such as impedance or admittance control. In these approaches, RL algorithms are employed to learn optimal values for parameters such as stiffness [374], damping, and inertia, enabling the robot to adapt its compliance to varying task requirements and contact conditions [206]. This data-driven adaptation helps maintain safety by minimizing excessive contact forces, preventing instability, and improving task performance.
@@
Recent research has demonstrated the effectiveness of RL-based parameter learning in a variety of contact-rich scenarios. For example, RL has been used to adjust impedance parameters online to ensure safe and robust assembly [285, 194, 79, 97], insertion [47, 256, 335], and surface interaction tasks [65]. Some works incorporate safety constraints directly into the RL framework, such as limiting the maximum allowable force or ensuring stability margins during learning [48, 399, 374]. Others combine RL with model-based safety filters [296] or control barrier functions [377] to guarantee constraint satisfaction throughout the learning and execution process [374].
@@
Large vision-language(-action) “robot foundation” models enhance safe execution by converting unstructured scene and instruction inputs into structured, verifiable manipulation plans, then constraining low‑level interaction.
High-level language and vision modules parse tasks and safety rules (e.g., force caps, no-touch regions, human proximity) and compile them into machine-checkable constraints; mid-level policies parameterize compliant skills (pose/force/impedance) rather than direct torques; low-level safety filters (CBF/reachability/MPC shields) project commands online to maintain invariance and limits [307].
VLA with safe constraints and tactile modality (e.g., “safeVLA”, “TLA”-style). Semantic policies are restricted via schema-conformant action spaces and pre-execution checks; code-as-policies and structured prompting inject explicit guardrails (caps, forbidden zones) before motion is issued [367, 307, 388, 368, 113, 353, 49].
Semantic-to-physical constraint grounding. VLMs / VLAs translate natural-language rules and scene understanding (affordances, fragile surfaces) into explicit force/region/speed constraints that downstream safety layers can enforce [370].
@@ 3.1.3 Provable Safety Methods
Robotic safety has long been researched in classical control before learning-based methods became prevalent [103, 133, 17, 104]. Rigorous mathematical tools were developed to provide provable guarantees of stability and safety [31, 30, 272, 145]. This section aims to systematically describe foundational methods for certifying safety, including lyapunov-based methods, control barrier functions (CBF), contraction metrics, reachability analysis and invariant sets. While originally developed for classical model-based control, many of these tools are increasingly integrated in modern learning-based frameworks [314, 60, 131, 313, 369, 316]. This section aims to establish the control theoretical foundations for understanding and developing safe learning methods for contact-rich robotic tasks.
@@
Lyapunov-based methods build on the concept of a Lyapunov function, a scalar energy-like function $V(\boldsymbol{x})$ that decreases along trajectories of a dynamical system. Formally, consider a system $\dot{\boldsymbol{x}}={f}(\boldsymbol{x},\boldsymbol{u})$ with state $\boldsymbol{x}\in\mathcal{X}$ and control input $\boldsymbol{u}\in\mathcal{U}$. A continuously differentiable function $V:\mathbb{R}^{n}\to\mathbb{R}_{\geq 0}$ is a Lyapunov function for a closed-loop system if it is positive definite ($V(\boldsymbol{0})=0$, $V(\boldsymbol{x})>0$ for $\boldsymbol{x}\neq\boldsymbol{0}$) and its time derivative is negative semi-definite ($\dot{V}(\boldsymbol{x})\leq 0\,\forall x$). For safety, Lyapunov analysis is often used to certify invariant sets. If we can identify a set of safe states $\mathcal{S}(e.g.\,\mathcal{S}=\{\boldsymbol{x}\mid V(\boldsymbol{x})\leq c\})$ and a Lyapunov-like function $V(\boldsymbol{x})$ that decreases or stays constant on $\mathcal{S}$, then any trajectory starting in $\mathcal{S}$ will remain in $\mathcal{S}$ for all time, making $\mathcal{S}$ a forward-invariant safe set [133].
@@
Lyapunov-based methods are well-suited to scenarios where the main objective is stability, for example, regulating contact force to a desired value. Nearly all classical robot controllers, e.g., PD controllers or compliance controllers, are derived with an implicit Lyapunov function to prove stability [165, 167, 168]. Passivity-based control is another Lyapunov-related approach. By designing the closed-loop system to be passive, one can guarantee stable interactions, which means the robot can safely interact with the environment because it will not generate energy that could cause instability [312, 77, 102, 60, 131].
@@
Despite their broad use, a major challenge for Lyapunov methods is finding a suitable Lyapunov function for complex nonlinear systems [133]. For high-dimensional robots with complex contact dynamics, hand-designing a $V(\boldsymbol{x})$ that decreases in all conditions, especially across contact/non-contact modes, is difficult. In hybrid systems, standard Lyapunov theory does not directly apply and one may need multiple Lyapunov functions [35, 86, 203]. Another limitation is that classical Lyapunov methods are primarily designed to certify stability or asymptotic convergence of the closed-loop system. Unless the Lyapunov sublevel sets are explicitly constructed to be forward invariant and contained in a prescribed safe set, Lyapunov stability alone does not guarantee that state or input constraints are respected along the entire transient [247, 160]. In particular, a Lyapunov-stable controller may drive a robot back to its desired equilibrium after a disturbance but still allow temporary violations of safe force or workspace limits [204, 333]. Therefore, Lyapunov-based designs are well-suited for ensuring long-term stability or boundedness, but typically need to be augmented with barrier-type conditions to enforce state constraints and safety [8, 31].
@@
Control barrier functions (CBFs) provide a formal, Lyapunov-like tool specifically for enforcing safety constraints. While Lyapunov functions certify the stability of an equilibrium, barrier functions certify forward invariance of a safe set, which prove the system will never enter certain unsafe regions [247, 7, 8]. The typical setup is as follows: Define a continuously differentiable scalar boundary function $h(\boldsymbol{x})$ such that the safe set $\mathcal{C}$ is $\{\boldsymbol{x}\mid h(\boldsymbol{x})\geq 0\}$ and the unsafe set is $\{\boldsymbol{x}\mid h(\boldsymbol{x})<0\}$. As an example, the boundary $h(\boldsymbol{x})=0$ can describe the contact force reaching a limit. A valid CBF satisfies a condition that ensures $h(\boldsymbol{x})$ cannot decrease into the negative unsafe region. One common form: There exists an extended class-$\mathcal{K}$ function $\alpha(\cdot)$ such that for all states on the boundary of $\mathcal{C}$, we can choose control inputs $\boldsymbol{u}$ to yield $\dot{h}(\boldsymbol{x},\boldsymbol{u})\geq-\alpha(h(\boldsymbol{x}))$. If this holds, then whenever $h(\boldsymbol{x})$ is zero or positive, its derivative cannot drive $h(\boldsymbol{x})$ across the zero boundary into the unsafe region. Geometrically, $h(\boldsymbol{\boldsymbol{x}})$ can be seen as a ”barrier” that goes to zero at the safety boundary and typically to negative values outside the safe set. The CBF condition forces the controlled dynamics to hover at or increase $h(\boldsymbol{x})$ at the boundary, never allowing a drop below zero [333].
@@
CBFs are especially powerful for real-time enforcement of safety constraints. Unlike open-loop potential field methods [20, 237], CBFs offer a guarantee of safety because the CBF condition is a hard constraint on the control input.
@@
Despite their strengths, CBFs come with several challenges. First, they require a reasonably accurate model of the system and environment. Model uncertainty or external disturbances can break the guarantee. There is active research on robust CBFs and adaptive CBFs to account for uncertainty, but guaranteeing safety under uncertainty often leads to conservative controllers [239, 40, 322]. Second, there can be conflicts between multiple CBFs or between a CBF and control goals. It is possible that no single controller can simultaneously satisfy all constraints and achieve the desired motion. In such cases, typically one should relax constraints [120, 37]. Third, the design of CBFs becomes significantly more involved for safety constraints with high relative degree, such as contact and force constraints that depend on higher-order derivatives of the state, which are ubiquitous in contact-rich robotic tasks [36, 319, 332].
@@
Contraction metrics offer a fundamentally different viewpoint than the Lyapunov-based and CBF methods that provide scalar certificates for stability or safety with respect to particular equilibria or safe sets. Rather than analyzing the behavior of trajectories relative to a fixed point or a boundary, contraction theory studies the evolution of the differential distance between arbitrary trajectories of a dynamical system. A system is said to be contracting if infinitesimal displacements between trajectories converge exponentially to zero under an appropriate Riemannian metric. When such a contraction metric exists, all trajectories within a contraction region converge exponentially toward one another, independent of their initial conditions. This yields a strong form of incremental stability: not only do trajectories converge to a specific equilibrium, but they converge toward a unique flow or invariant trajectory structure of the system. The contraction metric thus serves as the formal certificate ensuring this incremental convergence and generalizes classical Lyapunov analysis to trajectory-to-trajectory stability [182, 188, 273, 203].
@@
This condition means the distance between two nearby solutions measured by $\boldsymbol{M}$ decays at least at rate $\lambda$. One can interpret it as a generalized Lyapunov condition applied to the difference dynamics. If we consider the error $|\tilde{\boldsymbol{x}}|=\boldsymbol{x}_{1}-\boldsymbol{x}_{2}$ between two trajectories $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, contraction theory guarantees $|\tilde{\boldsymbol{x}}|\leq ke^{-\lambda t}|\tilde{\boldsymbol{x}}(\boldsymbol{0})|$ for some $k>0$. A remarkable property is that if a system is contracting, it has a unique equilibrium or invariant trajectory that is globally exponentially stable [188]. In control contexts, we often speak of control contraction metrics (CCMs). A CCM extends the problem to controlled systems. It provides a condition under which a feedback controller can be designed so that the closed-loop system is contracting. If a CCM exists, one can synthesize a controller that achieves exponential tracking of any feasible reference trajectory [203].
@@
Contraction metrics are particularly useful in scenarios requiring robust trajectory tracking and disturbance rejection. A major challenge of this approach is finding a contraction metric for a given system, which can be challenging as finding a Lyapunov function. For complicated robots, $\boldsymbol{M}(\boldsymbol{x})$ might need to be very complex or even non-existent in closed form [84, 85]. Another challenge is that contraction typically addresses stability properties but not explicit state constraints [271].
Reachability analysis and invariant sets approaches compute the set of states from which the system can remain safe indefinitely or the set of states that could lead to an unsafe condition. Techniques like Hamilton-Jacobi (HJ) reachability solve a PDE to find the backward reachable set of an unsafe target [216, 17]. Consider a system whose dynamics are described by an ordinary differential equation $\dot{\boldsymbol{x}}=\boldsymbol{f}(\boldsymbol{x}(t),\boldsymbol{u}(t))$ where $\boldsymbol{x}$ is the system’s states, $\boldsymbol{u}(t)$ is the control input. To compute the reachable set, a value function $V(\boldsymbol{x},t)$ is typically defined, representing whether the system can reach the target set from state $\boldsymbol{x}$ at time $t$. This value function satisfies the HJ equation:
@@
By solving this equation, the boundary of the reachable set can be obtained, allowing for the analysis of the system’s behavior under various control strategies [212, 52, 184].
The limitation is the curse of dimensionality. Solving HJ PDEs or exhaustive reachability is only feasible for simplified systems [61, 18, 80]. In safe learning contexts, reachability is often used as a runtime safety filter, where the reachable set boundary serves as a constraint. The filter minimally adjusts the nominal control input to satisfy the safety condition, thereby guaranteeing forward invariance of the safe set without strictly limiting the learning agent’s performance [183, 260, 257, 139, 287].
@@ 3.2 Contact-rich Tasks
Our categorization is application-independent, however, it is possible to categorize the tasks also w.r.t. the application areas such as industry, household and healthcare [310]. Many contact-rich tasks are demonstrated for household applications that cannot be automated, as opposed to industry. These include general household tasks such as wiping a table or a vase [158, 106], inserting a plug [251, 309], and opening a door [127, 231]; and kitchen tasks such as cutting vegetables [217, 233], stirring or scooping food [146, 208], and rolling pizza [269]. Industry applications usually focus on flexible manufacturing tasks such as inserting a gear [230, 194], nut threading [226], building furniture [172], and disassembly [242, 258]. Healthcare applications mostly include the surgery operations [245, 286] and physically assisting humans (Sec. 3.2.4).